OpenStudy (anonymous):

Find the distance between (1, 1) and (3, 4).

OpenStudy (anonymous):

sqrt (x2 - x1)^2 + (y2 - y1)^2

OpenStudy (hoblos):

d^2 = (3-1)^2 + (4-1)^2 = 4+9=13 d=sqrt13

OpenStudy (apoorvk):

this is done using the distance formula... for any two points (x,y) and (x',y') the distance between them is.. \[\sqrt{(x-x')^{2} + (y-y')^{2}}\] so here \[^{13}\] should be it.

OpenStudy (apoorvk):

i mean root of 13, ie \[\sqrt{13}\]